# SPO: Sequential Monte Carlo Policy Optimisation

### v2

@@ 4.3 M-step
After completing the E-step (which generates $q_{i}$), we proceed with the M-step, which optimises eq. 3 with respect to $\pi$, parameterised by $\theta$. By eliminating terms that are independent of $\pi$, we optimise the following objective, corresponding to a maximum a posteriori estimation with respect to the distribution $q_{i}$:
@@
This optimisation can be viewed as projecting the non-parametric policy $q_{i}$ back to the space of parameterisable policies $\Pi_{\theta}$, as performed in expert iteration style methods such as AlphaZero. $p(\theta)$ represents a prior over the parameter $\theta$. Previous works find that utilising a prior for $\theta$ to be close to the estimate from the previous iteration $\theta_{i}$ leads to stable training [1, 69] and so we optimise the following constrained objective:
@@
s.t. $\quad\mathbb{E}_{\mu_{q_{i}}(s)}\left[\text{KL}\left(\pi(a|s,\theta_{i}),\pi(a|s,\theta)\right)\right]<\epsilon.$
@@ 4.4 Policy Improvement
Previous approaches to using SMC for RL either do not perform iterative policy improvement using SMC [60], or lack policy improvement constraints needed for iterative improvement [41]. Provided SMC provides perfect estimates of the analytic target distribution at each step $i$, Expectation Maximisation algorithm will guarantee that successive iterations will result in monotonic improvements in our lower bound objective, with details outlined in section G.3 [50].
@@
In practice we are unlikely to generate perfect estimates of the target through sample based inference and leave derivations regarding the impact of this estimation on overall convergence for future work. We also draw the connection between our EM optimisation method and Mirror Descent Guided Policy Search (MD-GPS) [54]. Our objective can be viewed as a specific instance of MD-GPS (see section G.4). Depending on whether dynamics are linear or not optimising the EM objective can then be viewed either as exact or approximate mirror descent [9]. Monotonic improvement guarantees in MD-GPS follow from those of mirror descent.
@@ 5 Experiments
5.1 Experimental Setup
In order to ensure the robustness of our conclusions, we follow the evaluation methodology proposed by Agarwal et al. [2], see appendix H for further details. We include individual results in appendix C, along with additional analysis measuring the statistical significance of our results.
Baselines: Our model-free baselines include PPO [65], MPO [1] and V-MPO [69] for both continuous and discrete environments. For model-based algorithms, we compare performance to AlphaZero (including search improvements from MuZero [63]) and an SMC method introduced by Piché et al. [60]  [^2] We use MDQN [76], and SAC [33] for discrete and continuous environments, respectively. Note that in the discrete case, SMC-ENT using SAC was very unstable, so we chose MDQN as a better alternative. (which we refer to as SMC-ENT due to its use of maximum entropy RL to train the proposal and value function used within SMC). Our AlphaZero benchmark follows the official open-source implementation [^3] Open-source implementation available at mctx.. For continuous environments we baseline our results to Sampled MuZero [37], a modern extension to MuZero for large and/or continuous action spaces. We utilise a true environment model, aligning our implementation of Sampled MuZero more closely with AlphaZero. Our core experiments configure SPO with 16 particles and a horizon of 4, and AlphaZero with 64 simulations to equalise search budgets. For remaining parameters, see section E.1 and section D.3. Each environment and algorithm were evaluated with five random seeds. The complete codebase will be released upon publication.
@@ 5.2 Results
Policy Improvement: In fig. 2 we show empirical evidence that SPO outperforms all baseline methods across both discrete (left) and continuous (right) benchmarks for policy improvement. This conclusion also holds for the per-environment results reported in appendix C. SMC-ENT [60] is a strong SMC based algorithm, however, SPO outperforms it for both discrete and continuous environments providing strong evidence of the direct benefit of using SMC posterior estimates for policy improvement. We also highlight the importance regularising policy optimisation, which is practically achieved by solving for $\eta^{*}$ eq. 7, and is re-estimated at each iteration. In appendix B we ablate the impact of this adaptive method by comparing varying fixed temperature schedules.
We demonstrate that SPO performance is statistically significant compared to AlphaZero across the evaluated environments, see appendix C for detailed results. The substantial variation in AlphaZero performance across these environments underscores the difficulty in tuning it for diverse settings. For instance, while AlphaZero performs well on Boxoban and HalfCheetah, its performance drops significantly on other discrete and continuous problems. This inconsistency poses a major challenge for its applicability to real-world problems. In contrast, SPO performs consistently well across all environments (both discrete and continuous), highlighting its general applicability and robustness to various problem settings.
@@ 6 Conclusion
Our work provides several key contributions. First, we show that SPO is a powerful policy improvement operator, outperforming our model-based and model-free baselines. Additionally, SPO is shown to be competitive across both continuous and discrete domains, without additional environment-specific alterations. This illustrates the effectiveness of SPO as a generic and robust approach, in contrast to prior methods that require domain-specific enhancements. Furthermore, the parallelisable nature of SPO results in the efficient scaling behaviour of the search method. This allows SPO to achieve a significant boost in performance at inference time by scaling the search budget. Finally, we demonstrate that scaling SPO results in faster wall-clock inference time compared to previous work utilising tree-based search methods. The presented work culminates in a versatile, neural-guided, sample-based planning method that demonstrates superior performance and scalability over baselines.
Limitations & Future Work: Our work exclusively uses exact world models of the environment in order to isolate the effects of various planning methods on performance. Extending our research to include a learned world model would broaden the applicability of SPO to more complex problems that may not have a perfect simulator and are therefore unsuitable for direct planning. Additionally, we apply Sequential Monte Carlo (SMC) only in deterministic settings. Adapting our approach for stochastic environments presents challenges. Specifically, the importance weight update in SMC can lead to the selection of dynamics most advantageous to the agent, potentially fostering risk-seeking behaviour in stochastic settings. However, mitigation strategies exist, as discussed in Levine [42]. Finally, exploring alternative inference algorithms within our method could further improve the estimation of the target distribution.
@@ A.3.2 Network Architecture
In practice, we found the smaller networks used in the original Brax publication to limit overall performance. We instead use a $4$except for SMC-ENTR,for which we use 3 to stabilise performance.
layer feed forward network to represent both the value and policy network. Aside from the output layers, all layers are of size 256 and use SiLU non-linearities. We use the same architecture for all algorithms. Unlike Rubik’s Cube and Sokoban, we do not use a shared torso to learn embeddings.
@@ B.1 E-step KL constraint
The KL constraint the the E-step within the Expectation Maximisation framework is an important component of ensuring policy improvement due to the KL term in the ELBO. We investigate the impact of this adaptive temperature for both discrete and continuous environments. We compare SPO with an adaptive temperature updated every iteration to a variety of fixed temperatures.
@@
This ablation shows that SPO with an adaptive temperature is among the top performing hyperparameter settings across all environments. However we also note that it is possible to tune a temperature that works well when considering a wide range of temperatures. This is consistent with previous results in Peng et al. [57] that also find practically for specific problems a fixed temperature can be used, of course in practice having an algorithm that can learn this parameter itself is practically beneficial, removing hyperparameter tuning costs, since the appropriate temperature is likely problem dependant.
@@ C.1 Detailed Summary Results
In Figures 8 and 9, we present the absolute metrics following the recommendations by Agarwal et al. [2], Colas et al. [13] and Gorsane et al. [28]. These metrics evaluate the best set of network parameters identified during 20 evenly spaced training evaluation intervals. The absolute evaluations measure performance across 10 times the number of episodes periodically assessed during training, resulting in a total of 1280 episodes evaluated. We report point estimates and their 95% confidence intervals, which were calculated using stratified bootstrapping. Agarwal et al. [2] advocate for the Interquartile Mean (IQM) as a more robust and valid point estimate metric. Our results indicate that SPO surpasses all other algorithms in every point estimate. Notably, the IQM and mean point estimates demonstrate statistical significance.
@@ C.3 Individual Environment Results
All individual task results are presented in Figures C.3, C.3, and C.3. Specifically, we present the IQM of returns achieved during the evaluation intervals throughout training.
@@
(a) Humanoid.
(b) Sokoban.
(a) Rubik’s Cube 7 scrambles
@@ Appendix D SPO
$\mathcal{L}_{\pi}(\theta)=-\sum_{s,a\sim{\mathcal{D}}}q_{i}(s,a)\log\pi_{\theta_{i}}(a|s)$
@@ F.1 Overview of methods
EM algorithms differ on a small number of dimensions through which the algorithm can be understood. The first dimension, $\mathcal{G}$, is the optimization objective that is maximized in the E-step. This includes whether advantages, Q-values, or rewards are used, and with respect to which policy. We then consider whether a trust region constrained is used both in the E-step and M-step.
@@
Note that $T$ refers to an episode length. In additon we add two important dimensions of the $\mathcal{G}$ estimate (breadth and depth). Methods like MPO estimate the analytic solution directly using a Q-function. This can be leveraged to estimate the target distribution for a selection of $M$ actions, but without leveraging depth, so rewards and future states are not used to improve the estimates. In contrast, V-MPO leverages depth to form an n-step TD estimate, but only for one of the actions, leading to a relatively rough estimate of the analytic solution.
@@ G.1 E-step Analytic solution
This gives the analytic form for the optimal distribution $q$,
@@ G.2 E-step KL constraining temperature
We now derive the dual function to be minimised in order to obtain $\eta$ within the analytic solution for $q$ which is crucial for enforcing the KL constraint.
@@
Now that we have expressions for $-\frac{\eta+\gamma}{\eta}$ and $q$ we can form the dual function $g(\eta)$ by substituting these terms back into the original Lagrangian. After simplifying we can recover the following expression.
@@
G.3 Proof of Proposition 1: Monotone Improvement Guarantee
@@
G.4 Connection to Mirror Descent Guided Policy Search

### v3

@@ 4.3 M-step
After completing the E-step (which generates $q_{i}$), we proceed with the M-step, which optimises eq. 3 with respect to $\pi$, parametrised by $\theta$. By eliminating terms that are independent of $\pi$, we optimise the following objective, corresponding to a maximum a posteriori estimation with respect to the distribution $q_{i}$:
@@
This optimisation can be viewed as projecting the non-parametric policy $q_{i}$ back to the space of parametrisable policies $\Pi_{\theta}$, as performed in expert iteration style methods such as AlphaZero. $p(\theta)$ represents a prior over the parameter $\theta$. Previous works find that utilising a prior for $\theta$ to be close to the estimate from the previous iteration $\theta_{i}$ leads to stable training [1, 76]. Therefore we assume a gaussian prior over the current policy parameters, see section G.5 where we show utilising such a prior leads to the following constrained objective:
@@
s.t. $\quad\mathbb{E}_{\mu_{q_{i}}(s)}\left[\text{KL}\left(\pi(a|s,\theta_{i}),\pi(a|s,\theta)\right)\right]<\epsilon_{m}.$
@@ 4.4 Policy Improvement
Previous approaches to using SMC for RL either do not perform iterative policy improvement using SMC [65], or lack policy improvement constraints needed for iterative improvement [44]. Assuming that SMC provides perfect estimates of the analytic target distribution at each step $i$, Expectation Maximisation algorithm will guarantee that successive iterations will result in monotonic improvements in our lower bound objective, with details outlined in section G.4 [54].
@@
In practice we are unlikely to generate perfect estimates of the target through sample based inference and leave derivations regarding the impact of this estimation on overall convergence for future work. We also draw the connection between our EM optimisation method and Mirror Descent Guided Policy Search (MD-GPS) [58]. Our objective can be viewed as a specific instance of MD-GPS (see section G.6). Depending on whether dynamics are linear or not, optimising the EM objective can be viewed either as exact or approximate mirror descent [11]. Monotonic improvement guarantees in MD-GPS follow from those of mirror descent.
@@ 5 Experiments
5.1 Experimental Setup
In order to ensure the robustness of our conclusions, we follow the evaluation methodology proposed by Agarwal et al. [3], see appendix H for further details. This evaluation methodology groups performance across tasks within an environment suite, enabling clearer conclusions over the significance of learning algorithms as a whole. We include individual results in appendix C, along with additional analysis measuring the statistical significance of our results.
Baselines: Our model-free baselines include PPO [72], MPO [1] and V-MPO [76] for both continuous and discrete environments. For model-based algorithms, we compare performance to AlphaZero (including search improvements from MuZero [68]) and an SMC method introduced by Piché et al. [65]  [^2] We use MDQN [83], and SAC [37] for discrete and continuous environments, respectively. Note that in the discrete case, SMC-ENT using SAC was very unstable, so we chose MDQN as a better alternative. (which we refer to as SMC-ENT due to its use of maximum entropy RL to train the proposal and value function used within SMC). Our AlphaZero benchmark follows the official open-source implementation [^3] Open-source implementation available at mctx.. For continuous environments we baseline our results to Sampled MuZero [41], a modern extension to MuZero for large and/or continuous action spaces. We utilise a true environment model, aligning our implementation of Sampled MuZero more closely with AlphaZero. Our core experiments configure SPO with 16 particles and a horizon of 4, and AlphaZero with 64 simulations to equalise search budgets. For remaining parameters, see section E.1 and section D.3. Each environment and algorithm were evaluated with five random seeds.
@@ 5.2 Results
Policy Improvement: In fig. 2 we show empirical evidence that SPO outperforms all baseline methods across both discrete (left) and continuous (right) benchmarks for policy improvement. This conclusion also holds for the per-environment results reported in appendix C. SMC-ENT [65] is a strong SMC based algorithm, however, SPO outperforms it for both discrete and continuous environments providing strong evidence of the direct benefit of using SMC posterior estimates for policy improvement. We also highlight the importance regularising policy optimisation, which is practically achieved by solving for $\eta^{*}$ eq. 7, and is re-estimated at each iteration. In appendix B we ablate the impact of this adaptive method by comparing varying fixed temperature schedules. While a good temperature can be found through expensive grid search for each new environment we find the adaptive temperature ensures stable updates leading to strong learning performance across all environments, with minimal overhead to compute $\eta^{*}$.
We find that SPO performance is statistically significant compared to AlphaZero across the evaluated environments, see appendix C for detailed results. The substantial variation in AlphaZero performance across these environments underscores the difficulty in tuning it for diverse settings. For instance, while AlphaZero performs well on Boxoban and HalfCheetah, its performance drops significantly on other discrete and continuous problems. This inconsistency poses a major challenge for its applicability to real-world problems. In contrast, SPO performs consistently well across all environments (both discrete and continuous), highlighting its general applicability and robustness to various problem settings.
@@ 6 Conclusion
Our work provides several key contributions. First, we show that SPO is a powerful policy improvement operator, outperforming our model-based and model-free baselines. Additionally, SPO is shown to be competitive across both continuous and discrete domains, without additional environment-specific alterations. This illustrates the effectiveness of SPO as a generic and robust approach, in contrast to prior methods that require domain-specific enhancements. Furthermore, the parallelisable nature of SPO results in efficient scaling behaviour of search. This allows SPO to achieve a significant boost in performance at inference time by scaling the search budget. Finally, we demonstrate that scaling SPO results in faster wall-clock inference time compared to previous work utilising tree-based search methods. The presented work culminates in a versatile, neural-guided, sample-based planning method that demonstrates superior performance and scalability over baselines.
Limitations & Future Work: This work demonstrates the efficacy of combining probabilistic inference with reinforcement learning and amortisation. While our work considers a relatively standard form of Sequential Monte Carlo, future work could investigate making improvements to this inference method, in order to further improve the estimate of the target distribution, while maintaining the scalability benefits demonstrated. We also draw the connection to Active Inference [27] which has been explored in the context of agent behaviour in complex environments, where Monte Carlo Tree Search has been used to scale previous methods along with deep learning [25]. Leveraging Sequential Monte Carlo in such methods along with amortisation could be a promising direction of research to further scale such methods. Our work exclusively uses exact world models of the environment in order to isolate the effects of various planning methods on performance. Extending our research to include a learned world model would broaden the applicability of SPO to more complex problems that may not have a perfect simulator and are therefore unsuitable for direct planning. Additionally, we apply Sequential Monte Carlo (SMC) only in deterministic settings. Adapting our approach for stochastic environments presents challenges. Specifically, the importance weight update in SMC can lead to the selection of dynamics most advantageous to the agent, potentially fostering risk-seeking behaviour in stochastic settings. However, mitigation strategies exist, as discussed in Levine [45].
Acknowledgments and Disclosure of Funding
The authors would like to thank Teodora Pandeva for detailed discussions on SPO theory and for detailed revisions of the manuscript. Clément Bonnet for brainstorming discussions and feedback and David Kuric for revisions of the final manuscript. We also thank members of the InstaDeep team for their support in the preparation of the final manuscript. Lastly we thank the anonymous reviewers for comments and helpful discussions that helped improve the final version of the paper.
Research was supported with Cloud TPUs from Google’s TPU Research Cloud (TRC). Matthew Macfarlane is supported by the LIFT-project 019.011, which is partly financed by the Dutch Research Council (NWO).
@@ A.3.2 Network Architecture
In practice, we found the smaller networks used in the original Brax publication to limit overall performance. We instead use a 4 layer feed forward network444except for SMC-ENTR,for which we use 3 to stabilise performance.
to represent both the value and policy network. Aside from the output layers, all layers are of size 256 and use SiLU non-linearities. We use the same architecture for all algorithms. Unlike Rubik’s Cube and Sokoban, we do not use a shared torso to learn embeddings.
@@ B.1 E-step KL constraint
The KL constraint for the E-step within the Expectation Maximisation framework is an important component of ensuring policy improvement, due to the KL term in the ELBO. We investigate the impact of the adaptive temperature for both discrete and continuous environments. We compare SPO with an adaptive temperature updated every iteration to a variety of fixed temperatures.
@@
This ablation shows that SPO with an adaptive temperature is among the top performing hyperparameter settings across all environments. However we also note that it is possible to tune a temperature that works well when considering a wide range of temperatures. This is consistent with previous results in Peng et al. [62] that also find practically for specific problems a fixed temperature can be used. Of course in practice having an algorithm that can learn this parameter itself is practically beneficial, removing the need for costly hyperparameter tuning, since the appropriate temperature is likely problem dependant.
Subsequently, we evaluated whether the partial optimisation of the temperature parameter $\eta$ effectively maintained the desired KL divergence constraint and how different values of this constraint affected performance.
Figure 7: (a) The estimated KL divergence between the prior policy $\pi$ and the target policy $q$ generated by SMC for different values of $\epsilon$ during training on the Brax Ant task. (b) Evaluation performance during training for different values of $\epsilon$.
As shown in Figure 7(a), the KL divergence constraint is tightly maintained throughout training for different values of $\epsilon$. Figure 7(b) demonstrates that, for the Ant task, allowing a larger KL divergence between the prior and target policies ($q$ and $\pi$, respectively) leads to improved performance.
@@ B.2 E-step Q-value Optimisation
B.3 SMC Target Estimation Validation
Estimating the target distribution accurately is important for policy improvement in SPO. Therefore, we aim to validate the ability of SMC search to better approximate the true target distribution than a simple 1-step function evaluation as is done in algorithms such as MPO [1]. To do this, while the true target distribution is not directly accessible, we approximate it using an unbiased Monte Carlo oracle. We use a large computational budget of 1280 rollouts, to the end of the episode, for every state and every action. We evaluate the impact of varying planning horizons (depth) and particle counts on the KL divergence between the SMC-estimated target policy and the Monte Carlo oracle in the Sokoban environment.
Figure 9: Comparison of the KL divergence to large Monte Carlo simulation of target policy for different planning horizons and particle counts for Sokoban
As shown in Figure 9, increasing the number of particles and the planning depth reduces the KL divergence to the oracle, indicating improved estimation of the target distribution. This aligns with SMC theory [22], which suggests that more particles and deeper planning lead to better approximations. These results highlight the importance of leveraging both breadth (more particles) and depth (longer planning horizons) in SMC for accurate target estimation.
@@ C.1 Detailed Summary Results
In Figures 10 and 11, we present the absolute metrics following the recommendations by Agarwal et al. [3], Colas et al. [15] and Gorsane et al. [32]. These metrics evaluate the best set of network parameters identified during 20 evenly spaced training evaluation intervals. The absolute evaluations measure performance across 10 times the number of episodes periodically assessed during training, resulting in a total of 1280 episodes evaluated. We report point estimates and their 95% confidence intervals, which were calculated using stratified bootstrapping. Agarwal et al. [3] advocate for the Interquartile Mean (IQM) as a more robust and valid point estimate metric. Our results indicate that SPO surpasses all other algorithms in every point estimate. Notably, the IQM and mean point estimates demonstrate statistical significance.
@@ C.3 Individual Environment Results
All individual task results are presented in fig. 14. Specifically, we present the IQM of returns achieved during the evaluation intervals throughout training.
@@
(c) Humanoid.
(d) Sokoban.
(e) Rubik’s Cube 7 scrambles
Figure 14: Performance across different environments.
@@ Appendix D SPO
$\mathcal{L}_{\pi}(\theta)=-\sum_{s,a\sim{\mathcal{D}}}q_{i}(a|s)\log\pi_{\theta_{i}}(a|s)$
@@ D.3 Hyperparameters
D.4 Intuition for Temperature and Exploration:
The temperature parameter $\eta$ plays a crucial role in balancing exploration and exploitation during the search process in SPO. In this work, it is derived via first choosing the desired KL divergence $\epsilon$ between the target policy $q_{i}$ generated by search and the current policy $\pi_{i}$. If the temperature is too low (i.e., $\eta$ is small), the exponential weighting $\exp\left(A^{\bar{\pi}}(s,a)/\eta\right)$ becomes sharply peaked around the highest-advantage actions. This causes the importance weights to concentrate on a few particles, and during resampling, the particles can collapse onto a single root action. Such collapse reduces diversity in the search and renders the rest of the planning process ineffective. Conversely, if the temperature is too high (i.e., $\eta$ is large), the weighting becomes flatter, leading to excessive exploration. While exploration is necessary, too much can prevent the search from effectively focusing on promising paths. This work shows that deriving this value via a target KL leads to stable and strong performance across different environments.
@@ F.1 Overview of methods
EM algorithms differ on a small number of dimensions through which the algorithm can be understood. The first dimension, $\mathcal{G}$, is the optimization objective that is maximized in the E-step. This includes whether advantages, Q-values, or rewards are used, and with respect to which policy. We then consider whether a trust region constraint is used both in the E-step and M-step.
@@
Note that $T$ refers to an episode length. In additon we add two important dimensions of the $\mathcal{G}$ estimate (breadth and depth). Methods like MPO estimate the analytic solution directly using a Q-function. This can be leveraged to estimate the target distribution for a selection of $M$ actions, but without leveraging depth, so rewards and future states are not used to improve the estimates. In contrast, V-MPO leverages depth to form an n-step TD estimate, but only for one of the actions, leading to a relatively poor estimate of the analytic solution.
@@ G.1 E-step Analytic solution
The analytic form for the optimal distribution $q$ can then be calculated as:
@@ G.2 E-step KL constraining temperature
We now derive the dual function to be minimised in order to obtain $\eta$ within the analytic solution for $q$, which is crucial for enforcing the KL constraint.
@@
Now that we have expressions for $-\frac{\eta+\gamma}{\eta}$ and $q$ we can form the dual function $g(\eta)$ by substituting these terms back into the original Lagrangian. After simplifying we recover the following expression.
@@
G.3 E step: Constraint Optimisation
In optimising eq. 4, the first term $\mathbb{E}_{q(a|s)}\left[Q^{q}(s,a)\right]$ is dependent on the scale of reward in the environment. This can make it hard to choose $\alpha$, as the first and second terms are on arbitrary scales. Practically this can mean that for each new environment a new $\alpha$ should be used requiring costly hyperparameter tuning to find. As explored in previous work [1, 76] we choose to enforce a hard constraint, as opposed to a soft constraint. Instead of choosing $\alpha$ we choose $\epsilon$ which is the maximum KL divergence between $q$ and $\pi$ which is practically less sensitive to reward scale. We found that choosing an $\epsilon$ to be far easier, resulting in a value that generalised across all environments explored. This is important as there has been a trend in recent years to add hyperparameters to Reinforcement Learning algorithms [2] resulting in the need for costly hyperparameter sweeps for algorithms to work on each new problem they are applied to.
G.4 Proof of Proposition 1: Monotone Improvement Guarantee
@@
G.5 M-step objective
In Reinforcement Learning it can be beneficial to constrain policies from moving too far from the current policy, often leading to increased stability or performance of an algorithm [69, 72]. In this work we utilise a Gaussian prior around $\theta_{i}$, our current value of $\theta$, optimised from the previous iteration.
Therefore, $\theta\sim\mathcal{N}(\mu,\Sigma)$ where $\mu=\theta_{i}$ and $\Sigma^{-1}=\lambda F(\theta_{i})$, where $F$ is the Fisher information matrix. The prior term in eq. 11 can then be written as $\log p(\theta)=-\lambda(\theta-\theta_{i})^{T}F(\theta_{i})^{-1}(\theta-\theta_{i})+c$. The first term in this expression is the quadratic approximation of the KL divergence [69] while the second term $c$ is a term that does not depend on $\theta$ and therefore when optimising eq. 11 with respect to $\theta$, can be dropped. Using this approximation we rewrite eq. 11 as:
$\max_{\theta}\mathbb{E}_{s\sim\mu_{q}(s)}\left[\mathbb{E}_{a\sim q(a|s)}\left[\log\pi(a|s,\theta)\right]-\lambda\,\text{KL}\left(\pi(a|s,\theta_{i})\parallel\pi(a|s,\theta)\right)\right]$
Like the E-step, choosing $\lambda$ can be non-trivial so we convert it into a hard constraint optimisation eq. 12. Note that $\epsilon$ in eq. 12 is different from eq. 5. We note that considering a Gaussian prior in the M-step is not strictly required for SPO performance, however practically it adds stability to training.
G.6 Connection to Mirror Descent Guided Policy Search